At the beginning of 2014, Robotics Inc. acquired a manufacturing facility for $12.1 million. $9.1 million of the purchase price was allocated to the building. Depreciation for 2014 and 2015 was calculated using the straight-line method, a 20-year useful life, and a $1.1 million residual value. In 2016, the estimates of useful life and residual value were changed to 15 years and $510,000, respectively.

What is depreciation on the building for 2016? (Enter your answer in whole dollars.)
Answer #1
Depreciation for 2014 =($9,100,000-$1,100,000) / 20 Years =$400,000
Depreciation for 2015 =($9,100,000-$1,100,000) / 20 Years =$400,000
Book value as on Jan 1,2016 =$9,100,000 - ($400,000*2) =$8,300,000
Depreciation fo 2016 =($8,300,000 - $510,000) / 13Years =$599,231